Use linear substitution to solve the following first-order differential equation


𝑑𝑦/𝑑π‘₯=(2π‘₯+𝑦)/(2π‘₯+𝑦+1)



1
Expert's answer
2021-11-09T14:25:14-0500
"u=2x+y"

"\dfrac{du}{dx}=2+\dfrac{dy}{dx}"

Substitute


"\dfrac{du}{dx}-2=\dfrac{u}{u+1}"

"\dfrac{du}{dx}=\dfrac{3u+2}{u+1}"

"\dfrac{u+1}{3u+2}du=dx"

Integrate


"\int\dfrac{u+1}{3u+2}du=\int dx"

"\int\dfrac{u+1}{3u+2}du=\dfrac{1}{3}\int\dfrac{3u+2}{3u+2}du+\dfrac{1}{3}\int\dfrac{1}{3u+2}du"

"=\dfrac{1}{3}u+\dfrac{1}{9}\ln|(3u+2)|+C_1"

"\dfrac{1}{3}u+\dfrac{1}{9}\ln|(3u+2)|=x+\dfrac{1}{9}\ln C"

"3u+\ln|(3u+2)|=9x+\ln C"

"(3u+2)e^{3u}=Ce^{9x}"

Then


"(6x+3y+2)e^{6x+2y}=Ce^{9x}"

"(6x+3y+2)e^{2y}=Ce^{3x}"